### v1

# On Glashier’s Partition Theorem

Abstract.
Glashier’s theorem states that the number of partitions of $n$ into parts which repeat at most m1m-1 times is equal to the number of partitions of $n$ into parts which are not divisible by $m$. The $m=2$ case is Euler’s famous partition theorem. Recently, Andrews, Kumar, and Yee gave two new partition functions $C(n)$ and $D(n)$ related to Euler’s theorem. Lin and Zhang extended their result to Glashier’s theorem by generalizing $C(n)$. We generalize $D(n)$, prove an analogous partition identity for the $m=3$ case, and provide a new series equal to Glashier’s product both in the finite and infinite cases.
Key words and phrases:
Partitions, Partition Identities, Euler’s Theorem, Glashier’s Theorem, $q$-series.
2020 Mathematics Subject Classification:
05A17, 05A19, 11P81, 11P84.
1. Introduction
Euler’s celebrated partition theorem [1] can be stated as
(1.1) $A(n)=B(n)$
where $A(n)$ is the number of partitions of $n$ into distinct parts and $B(n)$ is the number of partitions of $n$ into odd parts.
In a recent work, Andrews, Kumar, and Yee [3] have provided two more elegant companion partition functions to $A(n)$ and $B(n)$. In particular, they have obtained the following partition identity.
Theorem $1.$
([3, Theorem $1.1$]) For any positive integer $n$, we have
(1.2) $A(n)=B(n)=C(n+1)=\dfrac{1}{2}D(n+1),$
where $C(n)$ is the number of partitions of $n$ with largest part even and parts not exceeding half of the largest part are distinct and $D(n)$ is the number of partitions of $n$ into non-negative parts wherein the smallest part appear exactly twice and no other parts are repeated.
Let $m\geq 2$ be an integer. Define $A_{m}(n)$ to be the number of partitions of n into parts repeating less than $m$ times and $B_{m}(n)$ to be the number of partitions of $n$ into parts not divisible by $m$. In 1883, Glashier [4] obtained the following partition theorem which is a natural extension of Euler’s theorem (1.1).
Theorem 1.2.
([4]) For any non-negative integer $n$, we have
(1.3) $A_{m}(n)=B_{m}(n).$
For any integer $0\leq j\leq m-1$, define $B^{(j)}_{m}(n)$ to be the number of partitions of $n$ into parts not divisible by $m$ and largest part congruent to $j$ modulo $m$. Clearly,
(1.4) $B_{m}(n)=\sum\limits_{j=1}^{m-1}B^{(j)}_{m}(n).$
In a recent work, Lin and Zhang [5] generalized Andrews, Kumar, and Yee’s $C(n)$ function to the case of $m$-regular partitions (partitions whose parts are not divisible by $m$). It is noteworthy to point out that the cardinality of the set $B^{\prime}_{m}(n)$ in their paper is the same as $B^{(m-1)}_{m}(n)$.
Define $C_{m}(n)$ to be the number of partitions of $n$ in which the largest part is divisible by $m$ (say mjmj) and all parts less than or equal to $j$ repeat less than $m$ times. Also, define $C_{m}(0):=1$.
Theorem 1.3.
([5, Theorem $1.5$]) For any positive integer $n$, we have
(1.5) $B^{(m-1)}_{m}(n)=C_{m}(n+1).$
Remark 1.
Lin and Zhang [5] provided both $q$-series (through generating functions) and bijective proofs of Theorem 1.3.
Let us now recall some notations from the theory of $q$-series which can be found in [1]. For any complex variable AA and non-negative integer NN, the conventional $q$-Pochhammer symbol is defined as
$(A)_{N}=(A;q)_{N}$ $:=\Bigg\{\begin{array}[]{lr}1\qquad\qquad\quad\,\,\,\text{when }N=0,\\ \prod\limits_{i=0}^{N-1}(1-Aq^{i})\,\,\,\text{otherwise, and}\end{array}$
$(A)_{\infty}=(A;q)_{\infty}$ $:=\lim\limits_{N\rightarrow\infty}(A)_{N}\,\,\text{for}\,\,|q|<1.$
Also, let $\zeta_{n}=e^{\frac{2\pi i}{n}}$ denote the primitive $n^{\text{th}}$ root of unity and $\omega=\zeta_{3}$ denote the primitive cubic root of unity here and throughout the remainder of the paper.
Now, define $D_{m}(n)$ to be the number of partitions of $n$ into non-negative parts where the smallest part occurs exactly $m$ times and all other parts repeat less than $m$ times.
Theorem 1.4.
For any non-negative integer $n$, we have
(1.6) $C_{m}(n)=\dfrac{D_{m}(n)+E_{m}(n)}{m},$
where $E_{m}(n)$ is such that the generating function of $E_{m}(n)$ is
(1.7) $\sum\limits_{n=0}^{\infty}E_{m}(n)q^{n}=\sum\limits_{n=0}^{\infty}q^{mn}(q^{n+1};q)_{\infty}\sum\limits_{j=1}^{m-1}(\zeta^{j}_{m}q^{n+1};q)_{\infty}.$
The $m=3$ case of (1.7) corresponds to the following identity
Theorem 1.5.
(1.8) $\varepsilon(q):=\sum\limits_{n=0}^{\infty}E_{3}(n)q^{n}=2-q-2q^{2}+\sum\limits_{n=2}^{\infty}(-1)^{n}\chi(n-1)q^{\binom{n+1}{2}+1},$
where $\chi(n)$ is defined as
$\chi(n):=\omega^{n}+\omega^{-n}=\Bigg\{\begin{array}[]{lr}2\qquad\,\,\,\text{if }3\mid n,\\ -1\quad\,\,\,\,\text{if }3\nmid n.\end{array}$
The complexity of the generating functions involved seemed to mitigate against an analog of the $C(n)=\frac{1}{2}D(n)$ portion of Theorem 1.1; however, surprisingly, a similar result exists for $m=3$ which clearly follows from (1.6) and (1.8).
Theorem 1.6.
Let $n$ be any positive integer that is not equal to a triangular number plus 11, then
(1.9) $C_{3}(n)=\frac{1}{3}D_{3}(n).$
Combining Theorems 1.2, 1.3, and 1.4, we then have the following partition identity which is a natural extension of Theorem $1.$
Theorem 1.7.
For any positive integer $n$, we have
(1.10) $A_{m}(n)=B_{m}(n)=\sum\limits_{k=1}^{m-2}B^{(k)}_{m}(n)+C_{m}(n+1)=\sum\limits_{k=1}^{m-2}B^{(k)}_{m}(n)+\dfrac{D_{m}(n+1)+E_{m}(n+1)}{m}.$
In view of the generating function of $B_{2}(n)$ in [3, $(2.1)$], we also obtain the following finite $q$-series identity.
Theorem 1.8.
For any positive integer NN, we have
(1.11) $1+\sum\limits_{n=1}^{N}\sum\limits_{j=1}^{m-1}\dfrac{q^{mn-j}}{\left(\prod\limits_{r=1}^{m-j}(q^{r};q^{m})_{n}\right)\cdot\left(\prod\limits_{r=m-j+1}^{m-1}(q^{r};q^{m})_{n-1}\right)}=\dfrac{(q^{m};q^{m})_{N}}{(q;q)_{mN}}.$
Letting $N\rightarrow\infty$ in (1.11) gives us the following $q$-series identity.
Corollary 1.9.
(1.12) $1+\sum\limits_{n=1}^{\infty}\sum\limits_{j=1}^{m-1}\dfrac{q^{mn-j}}{\left(\prod\limits_{r=1}^{m-j}(q^{r};q^{m})_{n}\right)\cdot\left(\prod\limits_{r=m-j+1}^{m-1}(q^{r};q^{m})_{n-1}\right)}=\dfrac{(q^{m};q^{m})_{\infty}}{(q;q)_{\infty}}.$
Remark 2.
The $m=2$ case of (1.12) above was recently proved by Andrews and Bachraoui in [2, Lemma 11, (1.10)]. Clearly, the series on the left-hand side of (1.12) is the generating function of $\sum\limits_{j=1}^{m-1}B^{(j)}_{m}(n)$ and the product on the right-hand side of (1.12) is the generating function of $B_{m}(n)$. The same observation (restricting all parts to be at most mNmN) proves Theorem 1.8.
The rest of the paper is organized as follows. In Sections 2 and 3, we provide $q$-series (generating function) proofs of Theorem 1.4 and 1.6 respectively.
2. Proof of Theorem 1.4
From the definition of $C_{m}(n)$, we then have
$\sum\limits_{n=0}^{\infty}C_{m}(n)q^{n}$ $=\sum\limits_{n=0}^{\infty}\dfrac{\prod\limits_{j=1}^{n}(1+q^{j}+\ldots+q^{(m-1)j})q^{mn}}{(q^{n+1};q)_{(m-1)n}}$
$=\sum\limits_{n=0}^{\infty}\dfrac{(q^{m};q^{m})_{n}q^{mn}}{(q;q)_{mn}}$
(2.1) $=(q^{m};q^{m})_{\infty}\sum\limits_{n=0}^{\infty}\dfrac{q^{mn}}{(q;q)_{mn}(q^{mn+m};q^{m})_{\infty}}.$
Using Euler’s identity [1, p. 1919, ($2.2.5$)]
(2.2) $\dfrac{1}{(t;q)_{\infty}}=\sum\limits_{n=0}^{\infty}\dfrac{t^{n}}{(q;q)_{n}},$
(replacing $q$ by $q^{m}$ and letting $t=q^{mn+m}$) in (2.1), we get
$\sum\limits_{n=0}^{\infty}C_{m}(n)q^{n}$ $=(q^{m};q^{m})_{\infty}\sum\limits_{n=0}^{\infty}\dfrac{q^{mn}}{(q;q)_{mn}}\sum\limits_{j=0}^{\infty}\dfrac{q^{mnj+mj}}{(q^{m};q^{m})_{j}}$
$=(q^{m};q^{m})_{\infty}\sum\limits_{n,j=0}^{\infty}\dfrac{q^{mn+mnj+mj}}{(q;q)_{mn}(q^{m};q^{m})_{j}}$
$=(q^{m};q^{m})_{\infty}\sum\limits_{n,j=0}^{\infty}\dfrac{1}{m}\left(\sum\limits_{k=0}^{m-1}\zeta^{kn}_{m}\right)\dfrac{q^{n+nj+mj}}{(q;q)_{n}(q^{m};q^{m})_{j}}$
(2.3) $=\dfrac{(q^{m};q^{m})_{\infty}}{m}\sum\limits_{j=0}^{\infty}\dfrac{q^{mj}}{(q^{m};q^{m})_{j}}\sum\limits_{k=0}^{m-1}\sum\limits_{n=0}^{\infty}\dfrac{(\zeta^{k}_{m}q^{j+1})^{n}}{(q;q)_{n}}.$
Using (2.2) with $t=\zeta^{k}_{m}q^{j+1}$ in (2.3), we get
$\sum\limits_{n=0}^{\infty}C_{m}(n)q^{n}$ $=\dfrac{(q^{m};q^{m})_{\infty}}{m}\sum\limits_{j=0}^{\infty}\dfrac{q^{mj}}{(q^{m};q^{m})_{j}}\sum\limits_{k=0}^{m-1}\dfrac{1}{(\zeta^{k}_{m}q^{j+1};q)_{\infty}}$
$=\dfrac{(q^{m};q^{m})_{\infty}}{m}\sum\limits_{j=0}^{\infty}\dfrac{q^{mj}}{(q^{m};q^{m})_{j}(q^{j+1};q)_{\infty}}$
$\qquad\qquad+\dfrac{(q^{m};q^{m})_{\infty}}{m}\sum\limits_{j=0}^{\infty}\dfrac{q^{mj}}{(q^{m};q^{m})_{j}}\sum\limits_{k=1}^{m-1}\dfrac{1}{(\zeta^{k}_{m}q^{j+1};q)_{\infty}}$
$=\dfrac{1}{m}\sum\limits_{j=0}^{\infty}\dfrac{q^{mj}(q^{m(j+1)};q^{m})_{\infty}}{(q^{j+1};q)_{\infty}}$
$\qquad\qquad+\dfrac{(q^{m};q^{m})_{\infty}}{m}\sum\limits_{j=0}^{\infty}\dfrac{q^{mj}(q^{j+1};q)_{\infty}}{(q^{m};q^{m})_{j}(q^{m(j+1)};q^{m})_{\infty}}\sum\limits_{k=1}^{m-1}(\zeta^{k}_{m}q^{j+1};q)_{\infty}$
$=\dfrac{1}{m}\sum\limits_{j=0}^{\infty}q^{mj}\cdot\prod\limits_{i=m+1}^{\infty}(1+q^{i}+\ldots+q^{(m-1)i})$
$\qquad\qquad+\dfrac{1}{m}\sum\limits_{j=0}^{\infty}q^{mj}(q^{j+1};q)_{\infty}\sum\limits_{k=1}^{m-1}(\zeta^{k}_{m}q^{j+1};q)_{\infty}$
(2.4) $=\dfrac{1}{m}\sum\limits_{n=0}^{\infty}D_{m}(n)q^{n}+\dfrac{1}{m}\sum\limits_{n=0}^{\infty}E_{m}(n)q^{n},$
where the generating function of $D_{m}(n)$ clearly follows from its definition and the generating function of $E_{m}(n)$ follows from (1.7). Comparing coefficients of $q^{n}$ on both sides of (2.4) for $n>1$, we have
$C_{m}(n)=\dfrac{D_{m}(n)+E_{m}(n)}{m}$
which completes the proof of Theorem 1.4.
3. Proof of Theorem 1.5
From (1.7), it is clear that
(3.1) $\varepsilon(q)=(q;q)_{\infty}\sum\limits_{j=0}^{\infty}\dfrac{q^{3j}}{(q;q)_{j}}\left((\omega q^{j+1};q)_{\infty}+(\omega^{-1}q^{j+1};q)_{\infty}\right).$
Now, define $\chi(n)$ as follows
$\chi(n):=\omega^{n}+\omega^{-n}=\Bigg\{\begin{array}[]{lr}2\qquad\,\,\,\text{if }3\mid n,\\ -1\quad\,\,\,\,\text{if }3\nmid n.\end{array}$
Using [1, p. 1919, ($2.2.6$)] with $t=-\omega q^{j+1}$ and $t=-\omega^{-1}q^{j+1}$ in (3.1), we get
(3.2) $\varepsilon(q)=(q;q)_{\infty}\sum\limits_{j=0}^{\infty}\dfrac{q^{3j}}{(q;q)_{j}}\sum\limits_{n=0}^{\infty}\dfrac{(-1)^{n}\chi(n)q^{\binom{n+1}{2}+nj}}{(q;q)_{n}}.$
Using Euler’s identity [1, p. 1919, ($2.2.5$)] with $t=q^{n+3}$ after interchanging summations in (3.2), we then have
$\varepsilon(q)$ $=(q;q)_{\infty}\sum\limits_{n=0}^{\infty}\dfrac{(-1)^{n}\chi(n)q^{\binom{n+1}{2}}}{(q;q)_{n}(q^{n+3};q)_{\infty}}$
$=\sum\limits_{n=0}^{\infty}(-1)^{n}\chi(n)q^{\binom{n+1}{2}}(1-q^{n+1})(1-q^{n+2})$
$=\sum\limits_{n=0}^{\infty}(-1)^{n}\chi(n)q^{\binom{n+1}{2}}(1-q^{n+1}(1+q)+q^{(n+1)+(n+2)})$
$=\sum\limits_{n=0}^{\infty}(-1)^{n}\chi(n)q^{\binom{n+1}{2}}-(1+q)\sum\limits_{n=0}^{\infty}(-1)^{n}\chi(n)q^{\binom{n+2}{2}}+\sum\limits_{n=0}^{\infty}(-1)^{n}\chi(n)q^{\binom{n+3}{2}}$
$=\sum\limits_{n=0}^{\infty}(-1)^{n}\chi(n)q^{\binom{n+1}{2}}+(1+q)\sum\limits_{n=1}^{\infty}(-1)^{n}\chi(n-1)q^{\binom{n+1}{2}}$
$\qquad+\sum\limits_{n=2}^{\infty}(-1)^{n}\chi(n-2)q^{\binom{n+1}{2}}$
$=(2+q)+\sum\limits_{n=2}^{\infty}(-1)^{n}\chi(n)q^{\binom{n+1}{2}}-2q(1+q)+(1+q)\sum\limits_{n=2}^{\infty}(-1)^{n}\chi(n-1)q^{\binom{n+1}{2}}$
$\qquad+\sum\limits_{n=2}^{\infty}(-1)^{n}\chi(n-2)q^{\binom{n+1}{2}}$
$=2-q-2q^{2}+\sum\limits_{n=2}^{\infty}(-1)^{n}\left(\chi(n)+\chi(n-1)+\chi(n-2)\right)q^{\binom{n+1}{2}}$
$\qquad+\sum\limits_{n=2}^{\infty}(-1)^{n}\chi(n-1)q^{\binom{n+1}{2}+1}.$
Now, $\chi(n)+\chi(n-1)+\chi(n-2)=0$ for all $n$. Hence,
$\varepsilon(q)=2-q-2q^{2}+\sum\limits_{n=2}^{\infty}(-1)^{n}\chi(n-1)q^{\binom{n+1}{2}+1}$
which completes the proof of Theorem 1.5.
4. Concluding Remarks
It would be very interesting to provide a bijective proof of Theorem 1.6. The same procedure in the proof of Theorem 1.5 in Section 3 can be followed for $E_{m}(n)$ with $m\geq 4$; however, the outcome will be much messier.
Acknowledgments
The authors would like to thank Alexander Berkovich and Rahul Kumar for their kind interest.

### v2

@@
-On Glashier’s Partition Theorem
+On Glaisher’s Partition Theorem
 Abstract.
-Glashier’s theorem states that the number of partitions of $n$ into parts which repeat at most m1m-1 times is equal to the number of partitions of $n$ into parts which are not divisible by $m$.
+Glaisher’s theorem states that the number of partitions of $n$ into parts which repeat at most m1m-1 times is equal to the number of partitions of $n$ into parts which are not divisible by $m$.
 The $m=2$ case is Euler’s famous partition theorem.
 Recently, Andrews, Kumar, and Yee gave two new partition functions $C(n)$ and $D(n)$ related to Euler’s theorem.
-Lin and Zhang extended their result to Glashier’s theorem by generalizing $C(n)$.
-We generalize $D(n)$, prove an analogous partition identity for the $m=3$ case, and provide a new series equal to Glashier’s product both in the finite and infinite cases.
+Lin and Zhang extended their result to Glaisher’s theorem by generalizing $C(n)$.
+We generalize $D(n)$, prove an analogous partition identity for the $m=3$ case, and show that the general case is an example of an almost partition identity.
+We also provide a new series equal to Glaisher’s product both in the finite and infinite cases.
 Key words and phrases:
-Partitions, Partition Identities, Euler’s Theorem, Glashier’s Theorem, $q$-series.
+Partitions, Partition Identities, Euler’s Theorem, Glaisher’s Theorem, $q$-series, Almost Partition Identities.
 2020 Mathematics Subject Classification:
@@
 Theorem $1.$
-([3, Theorem $1.1$]) For any positive integer $n$, we have
+([4, Theorem $1.1$]) For any positive integer $n$, we have
 (1.2) $A(n)=B(n)=C(n+1)=\dfrac{1}{2}D(n+1),$
@@
 Define $A_{m}(n)$ to be the number of partitions of n into parts repeating less than $m$ times and $B_{m}(n)$ to be the number of partitions of $n$ into parts not divisible by $m$.
-In 1883, Glashier [4] obtained the following partition theorem which is a natural extension of Euler’s theorem (1.1).
+In 1883, Glaisher [5] obtained the following partition theorem which is a natural extension of Euler’s theorem (1.1).
 Theorem 1.2.
@@
 Theorem 1.3.
-([5, Theorem $1.5$]) For any positive integer $n$, we have
+([6, Theorem $1.5$]) For any positive integer $n$, we have
 (1.5) $B^{(m-1)}_{m}(n)=C_{m}(n+1).$
@@
 $(A)_{\infty}=(A;q)_{\infty}$ $:=\lim\limits_{N\rightarrow\infty}(A)_{N}\,\,\text{for}\,\,|q|<1.$
+Next, for non-negative integers $a$, $b$, we define the $q$-binomial coefficient as
+$\left[\begin{matrix}a+b\\ b\end{matrix}\right]_{q}:=\Bigg\{\begin{array}[]{lr}\dfrac{(q)_{a+b}}{(q)_{a}(q)_{b}}\quad\text{for }a,b\geq 0,\\ 0\qquad\qquad\text{otherwise}.\end{array}$
 Also, let $\zeta_{n}=e^{\frac{2\pi i}{n}}$ denote the primitive $n^{\text{th}}$ root of unity and $\omega=\zeta_{3}$ denote the primitive cubic root of unity here and throughout the remainder of the paper.
@@
 where $E_{m}(n)$ is such that the generating function of $E_{m}(n)$ is
-(1.7) $\sum\limits_{n=0}^{\infty}E_{m}(n)q^{n}=\sum\limits_{n=0}^{\infty}q^{mn}(q^{n+1};q)_{\infty}\sum\limits_{j=1}^{m-1}(\zeta^{j}_{m}q^{n+1};q)_{\infty}.$
+(1.7) $\varepsilon_{m}(q):=\sum\limits_{n=0}^{\infty}E_{m}(n)q^{n}=\sum\limits_{n=0}^{\infty}q^{mn}(q^{n+1};q)_{\infty}\sum\limits_{j=1}^{m-1}(\zeta^{j}_{m}q^{n+1};q)_{\infty}.$
 The $m=3$ case of (1.7) corresponds to the following identity
 Theorem 1.5.
-(1.8) $\varepsilon(q):=\sum\limits_{n=0}^{\infty}E_{3}(n)q^{n}=2-q-2q^{2}+\sum\limits_{n=2}^{\infty}(-1)^{n}\chi(n-1)q^{\binom{n+1}{2}+1},$
-where $\chi(n)$ is defined as
-$\chi(n):=\omega^{n}+\omega^{-n}=\Bigg\{\begin{array}[]{lr}2\qquad\,\,\,\text{if }3\mid n,\\ -1\quad\,\,\,\,\text{if }3\nmid n.\end{array}$
+(1.8) $\varepsilon_{3}(q):=\sum\limits_{n=0}^{\infty}E_{3}(n)q^{n}=2-q-2q^{2}+\sum\limits_{n=2}^{\infty}(-1)^{n}\chi_{3}(n-1)q^{\binom{n+1}{2}+1},$
+where $\chi_{3}(n)$ is defined as
+$\chi_{3}(n):=\omega^{n}+\omega^{-n}=\Bigg\{\begin{array}[]{lr}2\qquad\,\,\,\text{if }3\mid n,\\ -1\quad\,\,\,\,\text{if }3\nmid n.\end{array}$
 The complexity of the generating functions involved seemed to mitigate against an analog of the $C(n)=\frac{1}{2}D(n)$ portion of Theorem 1.1; however, surprisingly, a similar result exists for $m=3$ which clearly follows from (1.6) and (1.8).
@@
 (1.9) $C_{3}(n)=\frac{1}{3}D_{3}(n).$
+Even more strikingly, we obtain another example of an almost partition identity (they are true for almost all $n$ but fail for an infinity of values; see [3]) for the general case which is as follows.
+Theorem 1.7.
+The identity
+(1.10) $C_{m}(n)=\frac{1}{m}D_{m}(n)$
+is true for almost all $n$. Moreover, if $N(x)$ is the number of times that (1.10) is true for $n<x$, then
+$\lim\limits_{x\rightarrow\infty}\frac{N(x)}{x}=1.$
 Combining Theorems 1.2, 1.3, and 1.4, we then have the following partition identity which is a natural extension of Theorem $1.$
@@
 (1.11) $A_{m}(n)=B_{m}(n)=\sum\limits_{k=1}^{m-2}B^{(k)}_{m}(n)+C_{m}(n+1)=\sum\limits_{k=1}^{m-2}B^{(k)}_{m}(n)+\dfrac{D_{m}(n+1)+E_{m}(n+1)}{m}.$
-In view of the generating function of $B_{2}(n)$ in [3, $(2.1)$], we also obtain the following finite $q$-series identity.
+In view of the generating function of $B_{2}(n)$ in [4, $(2.1)$], we also obtain the following finite $q$-series identity.
 Theorem 1.9.
@@
 The same observation (restricting all parts to be at most mNmN) proves Theorem 1.9.
-The rest of the paper is organized as follows. In Sections 2 and 3, we provide $q$-series (generating function) proofs of Theorem 1.4 and 1.6 respectively.
+The rest of the paper is organized as follows. In Sections 2, 3, and 4, we provide $q$-series (generating function) proofs of Theorem 1.4, 1.5, and 1.7 respectively.
 2. Proof of Theorem 1.4
@@
 From (1.7), it is clear that
-(3.1) $\varepsilon(q)=(q;q)_{\infty}\sum\limits_{j=0}^{\infty}\dfrac{q^{3j}}{(q;q)_{j}}\left((\omega q^{j+1};q)_{\infty}+(\omega^{-1}q^{j+1};q)_{\infty}\right).$
-Now, define $\chi(n)$ as follows
-$\chi(n):=\omega^{n}+\omega^{-n}=\Bigg\{\begin{array}[]{lr}2\qquad\,\,\,\text{if }3\mid n,\\ -1\quad\,\,\,\,\text{if }3\nmid n.\end{array}$
+(3.1) $\varepsilon_{3}(q)=(q;q)_{\infty}\sum\limits_{j=0}^{\infty}\dfrac{q^{3j}}{(q;q)_{j}}\left((\omega q^{j+1};q)_{\infty}+(\omega^{-1}q^{j+1};q)_{\infty}\right).$
+Now, define $\chi_{3}(n)$ as follows
+$\chi_{3}(n):=\omega^{n}+\omega^{-n}=\Bigg\{\begin{array}[]{lr}2\qquad\,\,\,\text{if }3\mid n,\\ -1\quad\,\,\,\,\text{if }3\nmid n.\end{array}$
 Using [1, p. 1919, ($2.2.6$)] with $t=-\omega q^{j+1}$ and $t=-\omega^{-1}q^{j+1}$ in (3.1), we get
-(3.2) $\varepsilon(q)=(q;q)_{\infty}\sum\limits_{j=0}^{\infty}\dfrac{q^{3j}}{(q;q)_{j}}\sum\limits_{n=0}^{\infty}\dfrac{(-1)^{n}\chi(n)q^{\binom{n+1}{2}+nj}}{(q;q)_{n}}.$
+(3.2) $\varepsilon_{3}(q)=(q;q)_{\infty}\sum\limits_{j=0}^{\infty}\dfrac{q^{3j}}{(q;q)_{j}}\sum\limits_{n=0}^{\infty}\dfrac{(-1)^{n}\chi_{3}(n)q^{\binom{n+1}{2}+nj}}{(q;q)_{n}}.$
 Using Euler’s identity [1, p. 1919, ($2.2.5$)] with $t=q^{n+3}$ after interchanging summations in (3.2), we then have
-$\varepsilon(q)$ $=(q;q)_{\infty}\sum\limits_{n=0}^{\infty}\dfrac{(-1)^{n}\chi(n)q^{\binom{n+1}{2}}}{(q;q)_{n}(q^{n+3};q)_{\infty}}$
-$=\sum\limits_{n=0}^{\infty}(-1)^{n}\chi(n)q^{\binom{n+1}{2}}(1-q^{n+1})(1-q^{n+2})$
-$=\sum\limits_{n=0}^{\infty}(-1)^{n}\chi(n)q^{\binom{n+1}{2}}(1-q^{n+1}(1+q)+q^{(n+1)+(n+2)})$
-$=\sum\limits_{n=0}^{\infty}(-1)^{n}\chi(n)q^{\binom{n+1}{2}}-(1+q)\sum\limits_{n=0}^{\infty}(-1)^{n}\chi(n)q^{\binom{n+2}{2}}+\sum\limits_{n=0}^{\infty}(-1)^{n}\chi(n)q^{\binom{n+3}{2}}$
-$=\sum\limits_{n=0}^{\infty}(-1)^{n}\chi(n)q^{\binom{n+1}{2}}+(1+q)\sum\limits_{n=1}^{\infty}(-1)^{n}\chi(n-1)q^{\binom{n+1}{2}}$
-$\qquad+\sum\limits_{n=2}^{\infty}(-1)^{n}\chi(n-2)q^{\binom{n+1}{2}}$
-$=(2+q)+\sum\limits_{n=2}^{\infty}(-1)^{n}\chi(n)q^{\binom{n+1}{2}}-2q(1+q)+(1+q)\sum\limits_{n=2}^{\infty}(-1)^{n}\chi(n-1)q^{\binom{n+1}{2}}$
-$\qquad+\sum\limits_{n=2}^{\infty}(-1)^{n}\chi(n-2)q^{\binom{n+1}{2}}$
-$=2-q-2q^{2}+\sum\limits_{n=2}^{\infty}(-1)^{n}\left(\chi(n)+\chi(n-1)+\chi(n-2)\right)q^{\binom{n+1}{2}}$
-$\qquad+\sum\limits_{n=2}^{\infty}(-1)^{n}\chi(n-1)q^{\binom{n+1}{2}+1}.$
-Now, $\chi(n)+\chi(n-1)+\chi(n-2)=0$ for all $n$. Hence,
-$\varepsilon(q)=2-q-2q^{2}+\sum\limits_{n=2}^{\infty}(-1)^{n}\chi(n-1)q^{\binom{n+1}{2}+1}$
+$\varepsilon_{3}(q)$ $=(q;q)_{\infty}\sum\limits_{n=0}^{\infty}\dfrac{(-1)^{n}\chi_{3}(n)q^{\binom{n+1}{2}}}{(q;q)_{n}(q^{n+3};q)_{\infty}}$
+$=\sum\limits_{n=0}^{\infty}(-1)^{n}\chi_{3}(n)q^{\binom{n+1}{2}}(1-q^{n+1})(1-q^{n+2})$
+$=\sum\limits_{n=0}^{\infty}(-1)^{n}\chi_{3}(n)q^{\binom{n+1}{2}}(1-q^{n+1}(1+q)+q^{(n+1)+(n+2)})$
+$=\sum\limits_{n=0}^{\infty}(-1)^{n}\chi_{3}(n)q^{\binom{n+1}{2}}-(1+q)\sum\limits_{n=0}^{\infty}(-1)^{n}\chi_{3}(n)q^{\binom{n+2}{2}}+\sum\limits_{n=0}^{\infty}(-1)^{n}\chi_{3}(n)q^{\binom{n+3}{2}}$
+$=\sum\limits_{n=0}^{\infty}(-1)^{n}\chi_{3}(n)q^{\binom{n+1}{2}}+(1+q)\sum\limits_{n=1}^{\infty}(-1)^{n}\chi_{3}(n-1)q^{\binom{n+1}{2}}$
+$\qquad+\sum\limits_{n=2}^{\infty}(-1)^{n}\chi_{3}(n-2)q^{\binom{n+1}{2}}$
+$=(2+q)+\sum\limits_{n=2}^{\infty}(-1)^{n}\chi_{3}(n)q^{\binom{n+1}{2}}-2q(1+q)+(1+q)\sum\limits_{n=2}^{\infty}(-1)^{n}\chi_{3}(n-1)q^{\binom{n+1}{2}}$
+$\qquad+\sum\limits_{n=2}^{\infty}(-1)^{n}\chi_{3}(n-2)q^{\binom{n+1}{2}}$
+$=2-q-2q^{2}+\sum\limits_{n=2}^{\infty}(-1)^{n}\left(\chi_{3}(n)+\chi_{3}(n-1)+\chi_{3}(n-2)\right)q^{\binom{n+1}{2}}$
+$\qquad+\sum\limits_{n=2}^{\infty}(-1)^{n}\chi_{3}(n-1)q^{\binom{n+1}{2}+1}.$
+Now, $\chi_{3}(n)+\chi_{3}(n-1)+\chi_{3}(n-2)=0$ for all $n$. Hence,
+$\varepsilon_{3}(q)=2-q-2q^{2}+\sum\limits_{n=2}^{\infty}(-1)^{n}\chi_{3}(n-1)q^{\binom{n+1}{2}+1}$
 which completes the proof of Theorem 1.5.
+4. Proof of Theorem 1.7
+Define $\chi_{m}(n)$ as follows
+$\chi_{m}(n):=\sum\limits_{j=1}^{m-1}\zeta^{jn}_{m}=\Bigg\{\begin{array}[]{lr}m-1\,\,\,\text{if }m\mid n,\\ -1\quad\,\,\,\,\,\text{if }m\nmid n.\end{array}$
+Using [1, p. 1919, ($2.2.6$)] with $t=-\zeta^{j}_{m}q^{n}$ in (1.7), we get
+(4.1) $\varepsilon_{m}(q)$ $=(q;q)_{\infty}\sum\limits_{n=0}^{\infty}\dfrac{q^{mn}}{(q)_{n}}\sum\limits_{k=0}^{\infty}\dfrac{(-1)^{k}\chi_{m}(k)q^{\binom{k+1}{2}+nk}}{(q;q)_{k}}.$
+Using Euler’s identity [1, p. 1919, ($2.2.5$)] with $t=q^{k+m}$ after interchanging summations in (4.1), we then have
+$\varepsilon_{m}(q)$ $=(q;q)_{\infty}\sum\limits_{k=0}^{\infty}\dfrac{(-1)^{k}\chi_{m}(k)q^{\binom{k+1}{2}}}{(q;q)_{k}(q^{k+m};q)_{\infty}}$
+(4.2) $=\sum\limits_{k=0}^{\infty}(-1)^{k}\chi_{m}(k)q^{\binom{k+1}{2}}(q^{k+1};q)_{m-1}.$
+Using [1, p. $36$, ($3.3.6$)] with the substitutions N=m1N=m-1 and $z=q^{k+1}$, we get
+$\varepsilon_{m}(q)$ $=\sum\limits_{k=0}^{\infty}(-1)^{k}\chi_{m}(k)q^{\binom{k+1}{2}}\sum\limits_{j=0}^{m-1}(-1)^{j}\left[\begin{matrix}m-1\\ j\end{matrix}\right]_{q}q^{\binom{j+1}{2}+kj}$
+$=\sum\limits_{k=0}^{\infty}(-1)^{k}\chi_{m}(k)\sum\limits_{j=0}^{m-1}(-1)^{j}\left[\begin{matrix}m-1\\ j\end{matrix}\right]_{q}q^{\binom{j+k+1}{2}}$
+$=\sum\limits_{j=0}^{m-1}(-1)^{j}\left[\begin{matrix}m-1\\ j\end{matrix}\right]_{q}\sum\limits_{k=0}^{\infty}(-1)^{k}\chi_{m}(k)q^{\binom{j+k+1}{2}}$
+$=\sum\limits_{j=0}^{m-1}\left[\begin{matrix}m-1\\ j\end{matrix}\right]_{q}\sum\limits_{k=j}^{\infty}(-1)^{k}\chi_{m}(k-j)q^{\binom{k+1}{2}}$
+$=-\sum\limits_{j=0}^{m-1}\left[\begin{matrix}m-1\\ j\end{matrix}\right]_{q}\sum\limits_{k=0}^{j-1}(-1)^{k}\chi_{m}(k-j)q^{\binom{k+1}{2}}$
+$\qquad+\sum\limits_{j=0}^{m-1}\left[\begin{matrix}m-1\\ j\end{matrix}\right]_{q}\sum\limits_{k=0}^{\infty}(-1)^{k}\chi_{m}(k-j)q^{\binom{k+1}{2}}$
+$=-\sum\limits_{j=0}^{m-1}\left[\begin{matrix}m-1\\ j\end{matrix}\right]_{q}\sum\limits_{k=0}^{j-1}(-1)^{k}\chi_{m}(k-j)q^{\binom{k+1}{2}}$
+$\qquad+\sum\limits_{k=0}^{\infty}(-1)^{k}q^{\binom{k+1}{2}}\sum\limits_{j=0}^{m-1}\chi_{m}(k-j)\left[\begin{matrix}m-1\\ j\end{matrix}\right]_{q}$
+$=-\sum\limits_{j=0}^{m-1}\left[\begin{matrix}m-1\\ j\end{matrix}\right]_{q}\sum\limits_{k=0}^{j-1}(-1)^{k}\chi_{m}(k-j)q^{\binom{k+1}{2}}$
+$\qquad+\sum\limits_{k=0}^{\infty}(-1)^{k}q^{\binom{k+1}{2}}\sum\limits_{j=0}^{m-1}\chi_{m}(k-j)\left(\left[\begin{matrix}m-1\\ j\end{matrix}\right]_{q}-1\right),$
+where we have used the fact that $\sum\limits_{j=0}^{m-1}\chi_{m}(k-j)=0$ in the second sum on the last line above. Hence, we have
+(4.3) $\varepsilon_{m}(q)=P_{m}(q)+\sum\limits_{k=0}^{\infty}(-1)^{k}q^{\binom{k+1}{2}}\sum\limits_{j=0}^{m-1}\chi_{m}(k-j)\left(\left[\begin{matrix}m-1\\ j\end{matrix}\right]_{q}-1\right),$
+where $P_{m}(q):=-\sum\limits_{j=0}^{m-1}\left[\begin{matrix}m-1\\ j\end{matrix}\right]_{q}\sum\limits_{k=0}^{j-1}(-1)^{k}\chi_{m}(k-j)q^{\binom{k+1}{2}}$.
+The polynomial
+$\sum\limits_{j=0}^{m-1}\chi_{m}(k-j)\left(\left[\begin{matrix}m-1\\ j\end{matrix}\right]_{q}-1\right)$
+in (4.3) above has at most $2^{m-1}-m$ terms.
+Now, there are at most $\sqrt{2x}+1$ triangular numbers less than $x$.
+Hence, the maximum possible number of non-zero coefficients in $\varepsilon_{m}(q)$ among the first $x$ terms is at most $(2^{m-1}-m)\cdot(\sqrt{2x}+1)$ leaving the number of coefficients that vanish as less than
+$x-(2^{m-1}-m)\cdot(\sqrt{2x}+1)$
+which is asymptotic to $x$ as $x\rightarrow\infty$. This shows that $E_{m}(n)$ is almost always zero which completes the proof of Theorem 1.7.
 5. Concluding Remarks
 It would be very interesting to provide a bijective proof of Theorem 1.6.
-The same procedure in the proof of Theorem 1.5 in Section 3 can be followed for $E_{m}(n)$ with $m\geq 4$; however, the outcome will be much messier.
 Acknowledgments